### v2

@@ 3.1 Sequential repetition of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$
Analogously to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, in this section we will show that security for sequential repetition follows from the payoff function for entangled ($ent$) attackers $T^{ent}_{i}(\textbf{a}_{i})=\gamma_{\textsc{c}}\textbf{1}_{\textsc{c}}(\textbf{a}_{i})-\gamma_{\perp}\textbf{1}_{\perp}(\textbf{a}_{i})-\gamma_{\textsc{i}}\textbf{1}_{\textsc{i}}(\textbf{a}_{i})$, where $(\gamma_{\textsc{c}},\gamma_{\perp},\gamma_{\textsc{c}})=\frac{1}{\sqrt{488625947}}(943,1107,22057)$. Theorem 2.9 will be recovered for entangled attackers with $q\leq n/2-5$ for $\alpha^{ent}:=\gamma_{\textsc{c}}\eta(1-p_{err})-\gamma_{\perp}(1-\eta)-\gamma_{\textsc{i}}\eta p_{err}$.
Let $\mathcal{A}\subseteq\Delta_{2}$ be the set of probabilities $\textbf{q}=(q_{\textsc{c}},q_{\perp},q_{\textsc{i}})$ that the attackers can reproduce. Notice that since for $\textbf{q}_{1},\textbf{q}_{2}\in\mathcal{A}$, the attackers can play any convex combination of them, the set $\mathcal{A}$ is convex. Let $\Delta=0.013$ and, for simplicity, denote $\mathfrak{b}_{\eta}:=\frac{1}{4}\big{(}1-(\tilde{w}^{\xi}(\eta)+\Delta)$. Let $\mathcal{R}\subset\Delta_{2}$ be the region of the probability simplex defined as $\mathcal{R}:=\cup_{\eta\in(0.53,1]}\mathcal{R}_{\eta}\subset\Delta_{2}$, where the sets $\mathcal{R}_{\eta}$ are defined as
$\begin{split}\mathcal{R}_{\eta}=\big{\{}(p_{\textsc{c}},p_{\perp},p_{\textsc{i}})\in\Delta_{2}\mid p_{\textsc{c}}&>\max\{1-p_{\perp}+\frac{\mathfrak{b}_{\eta}(1-2p_{\perp})}{1-2(\eta-\xi)},\frac{\eta-\xi-\cos^{2}(\frac{\pi}{8})-\mathfrak{b}_{\eta}}{1-(\eta-\xi)}p_{\perp}+\cos^{2}(\frac{\pi}{8})\}\\ \wedge p_{\textsc{i}}&<\min\{\frac{\mathfrak{b}_{\eta}(1-2p_{\perp})}{2(\eta-\xi)-1},\frac{\mathfrak{b}_{\eta}-\sin^{2}(\frac{\pi}{8})}{1-(\eta-\xi)}p_{\perp}+\sin^{2}(\frac{\pi}{8})\}\big{\}}.\end{split}$ (44)
@@
Under the conditions of Theorem 3.12, any q that the attackers can reproduce is such that $\textbf{q}\notin\mathcal{R}$, i.e., $\mathcal{R}\cap\mathcal{A}=\emptyset$, and thus
@@
Let $\textbf{q}=(q_{\textsc{c}},q_{\perp},q_{\textsc{i}})$ be the probability distribution corresponding to an attack from Alice and Bob and define $\mathcal{A}$ as the set of all probabilities If they are allowed to answer with probability $\eta\pm\xi$, then
$\textbf{q}=(q_{\textsc{c}},1-(\eta+\lambda),q_{\textsc{i}}),\lambda\in[-\xi,\xi].$ (46)
@@
$q_{\textsc{i}}\geq\mathfrak{b}_{\eta},\forall\lambda\in[-\xi,\xi].$ (47)
@@
$\mathcal{R}^{0}_{\eta}=\{(p_{\textsc{c}},p_{\perp},p_{\textsc{i}})\in\Delta_{2}\mid p_{\perp}\in[1-(\eta-\xi),1-(\eta+\xi)]\wedge p_{\textsc{i}}<\mathfrak{b}_{\eta}\},$ (48)
@@
Notice that the strategies $S_{\textsc{tfkw}}$ and $S_{guess}$, reproduce $\textbf{q}_{\textsc{tfkw}}=(\cos^{2}(\frac{\pi}{8}),0,1-\cos^{2}(\frac{\pi}{8}))$ and $\textbf{q}_{guess}=(\frac{1}{2},\frac{1}{2},0)$ and, respectively, and therefore $\textbf{q}_{\textsc{tfkw}},\textbf{q}_{guess}\in\mathcal{A}$. The attackers can also do the strategies consisting on always responding $\perp$ or always being incorrect (e.g answering inconsistent answers $a\neq b$), which reproduce the probabilities $\textbf{q}=(0,1,0)$ and $\textbf{q}=(0,0,1)$, respectively. Therefore, the convex hull of these four points is a convex subset of $\mathcal{A}$. Consider the straight line $s_{1}$ defined by the two points $\textbf{q}_{guess}$ and $(\eta-\xi-\mathfrak{b}_{\eta},1-(\eta-\xi),\mathfrak{b}_{\eta})\in\partial\mathcal{R}^{0}_{\eta}$, where $\partial$ denotes the boundary, which is given by
$s1:\frac{x-\frac{1}{2}}{\eta-\xi-\frac{1}{2}-\mathfrak{b}_{\eta}}=\frac{y-\frac{1}{2}}{\frac{1}{2}-(\eta-\xi)}=\frac{z}{\mathfrak{b}_{\eta}}.$ (50)
Consider the set $\mathcal{R}^{1}_{\eta}$, defined as the points in $\Delta_{2}$ that do not belong to $\mathcal{R}^{0}_{\eta}$ that are ’below’ the straight line $s_{1}$ and are such that $p_{\perp}\leq 1-(\eta+\xi)$, formally defined as
$\begin{split}\mathcal{R}^{1}_{\eta}=\big{\{}(p_{\textsc{c}},p_{\perp},p_{\textsc{i}})\in\Delta_{2}\mid p_{\textsc{c}}>1-p_{\perp}+\frac{\mathfrak{b}_{\eta}(1-2p_{\perp})}{1-2(\eta-\xi)}\wedge p_{\textsc{i}}<\frac{\mathfrak{b}_{\eta}(1-2p_{\perp})}{2(\eta-\xi)-1}\big{\}}\setminus\mathcal{R}^{0}_{\eta}.\end{split}$ (51)
@@
Similarly, consider the set $\mathcal{R}^{2}_{\eta}$ consisting of the points in $\Delta_{2}$ that are such that $p_{\perp}\geq 1-(\eta+\xi)$ and that are ’below’ the straight line given by the points $\textbf{q}_{\textsc{tfkw}}$ and $(\eta-\xi-\mathfrak{b}_{\eta},1-(\eta-\xi),\mathfrak{b}_{\eta})$, formally defined as
$\begin{split}\mathcal{R}^{2}_{\eta}=\big{\{}(p_{\textsc{c}},p_{\perp},p_{\textsc{i}})\in\Delta_{2}\mid p_{\textsc{c}}>\frac{\eta-\xi-\cos^{2}(\frac{\pi}{8})-\mathfrak{b}_{\eta}}{1-(\eta-\xi)}p_{\perp}+\cos^{2}(\frac{\pi}{8})\}\wedge p_{\textsc{i}}<\frac{\mathfrak{b}_{\eta}-\sin^{2}(\frac{\pi}{8})}{1-(\eta-\xi)}p_{\perp}+\sin^{2}(\frac{\pi}{8})\big{\}}\end{split}$ (52)
@@
(Sequential repetition for entangled attackers) Let $\eta$ and $p_{err}$ be such that $\alpha^{ent}>0$. Then, any sequential strategy that attackers who pre-share $q\leq n/2-5$ qubits to break $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$  is such that $\mathbb{E}[\Gamma_{r}^{ent,att}]\leq 0$ and, moreover, the probability that they obtain a total score $r\alpha^{ent}$ that an honest party would is exponentially small:
$\mathbb{P}[\Gamma_{r}^{att}\geq r\alpha^{ent}]\leq e^{-r(\alpha^{ent})^{2}/2}.$ (53)
The values for which $\alpha^{ent}>0$ correspond to the points below the straight line in Fig. 3.
@@
Let $\textbf{p}_{1}=(959/1000,0,41/1000)$ and $\textbf{p}_{2}=(27/50,23/50,0)$. The line segment given by $\textbf{p}_{1}$ and $\textbf{p}_{2}$ is a subset of $\mathcal{R}$ and it can be described as the intersection of $\Delta_{2}$ and the plane $\gamma_{\textsc{c}}p_{\textsc{c}}-\gamma_{\perp}p\gamma_{\perp}-\gamma_{\textsc{i}}p_{\textsc{i}}=0$, with the normal vector $\bm{\gamma}=(\gamma_{\textsc{c}},\gamma_{\perp},\gamma_{\textsc{i}})$, where $\gamma_{\textsc{c}}=\frac{943}{\sqrt{488625947}}\simeq 0.04266$, $\gamma_{\perp}=\frac{1107}{\sqrt{488625947}}\simeq 0.050079$ and $\gamma_{\textsc{c}}=\frac{22057}{\sqrt{488625947}}\simeq 0.99783$. Consider the follow non-empty partition of the simplex $\Delta_{2}$:
@@ 4.1 Discrete uniform choice of basis over the Bloch sphere
We define one round of the $\mathrm{QPV}_{m_{\theta\varphi}}^{\eta}$ protocol as:
@@
    $V_{0}$ and $V_{1}$ secretly agree on $m_{\theta}$, $m_{\varphi}$ and $x\in[m]$. In addition, $V_{0}$ prepares the EPR pair $\left|\Omega\right\rangle=(\left|00\right\rangle+\left|11\right\rangle)/\sqrt{2}$.
@@
    $V_{0}$ sends one qubit $Q$ of $\left|\Omega\right\rangle$ to $P$ and $V_{1}$ sends $x$ to $P$ coordinating their times so that $Q$ and $x$ arrive at $P$ at the same time.
@@
    $P$ measures $Q$ in the basis $\{\left|0_{x}\right\rangle,\left|1_{x}\right\rangle\}$ and broadcasts her outcome to $V_{0}$ and $V_{1}$. If, due to experimental losses, $Q$ does not arrive at $P$, she broadcasts ‘no detection’ with the symbol $\perp$.
@@
    If $V_{0}$ and $V_{1}$ receive both the same bit at the corresponding time, they accept. If they receive $\perp$, they record ‘no photon’.
At the end of the protocol, when sequentially run $r$ times, the verifiers check that the relative frequency of $\perp$ is upper bounded by $1-\eta$. Notice that $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ is recovered for $m=2$, with the unique choice of $m_{\theta}=2$ and $m_{\varphi}=1$.
4.2 Security of the $\mathrm{QPV}_{m_{\theta\varphi}}^{\eta}$ protocol under photon loss and non-entangled attackers
In an analogy to the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol, an attack on the $\mathrm{QPV}_{m_{\theta\varphi}}^{\eta}$ protocol can be associated with a MoE (monogamy-of-entanglement) game in the following way. Let $V$ be the register of the qubit of the verifier, with associated Hilbert space $\mathcal{H}_{V}=\mathbb{C}^{2}$, with $\mathcal{X}=[m]$ and $\mathcal{V}=\{0,1\}$. The verifiers perform the collection of measurements

### v3

@@ 3.2 Sequential repetition of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$
We will next show that, under the conditions of Theorem 3.11, attackers will reproduce a score $\Gamma^{ent,att}_{r}$ that will fail the $\mathsf{T}^{rf-\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ with exponentially high probability. Let $\mathcal{A}\subseteq\Delta_{2}$ be the set of probabilities $\textbf{q}=(q_{\textsc{c}},q_{\perp},q_{\textsc{i}})$ that the attackers can reproduce, where $q_{\textsc{i}}:=q_{\textsc{w}}+q_{\lightning}$. Notice that since for $\textbf{q}_{1},\textbf{q}_{2}\in\mathcal{A}$, the attackers can play any convex combination of them, the set $\mathcal{A}$ is convex. Let $\Delta=0.013$ and, for simplicity, denote $\mathfrak{b}_{\eta}:=\frac{1}{4}\big{(}1-(\tilde{w}^{\epsilon}(\eta)+\Delta)$. Let $\mathcal{R}\subset\Delta_{2}$ be the region of the probability simplex defined as $\mathcal{R}:=\cup_{\eta\in(0.53,1]}\mathcal{R}_{\eta}\subset\Delta_{2}$, where the sets $\mathcal{R}_{\eta}$ are defined as
$\begin{split}\mathcal{R}_{\eta}=\big{\{}(p_{\textsc{c}},p_{\perp},p_{\textsc{i}})\in\Delta_{2}\mid p_{\textsc{c}}&>\max\{1-p_{\perp}+\frac{\mathfrak{b}_{\eta}(1-2p_{\perp})}{1-2(\eta-\epsilon)},\frac{\eta-\epsilon-\cos^{2}(\frac{\pi}{8})-\mathfrak{b}_{\eta}}{1-(\eta-\epsilon)}p_{\perp}+\cos^{2}(\frac{\pi}{8})\}\\ \wedge p_{\textsc{i}}&<\min\{\frac{\mathfrak{b}_{\eta}(1-2p_{\perp})}{2(\eta-\epsilon)-1},\frac{\mathfrak{b}_{\eta}-\sin^{2}(\frac{\pi}{8})}{1-(\eta-\epsilon)}p_{\perp}+\sin^{2}(\frac{\pi}{8})\}\big{\}}.\end{split}$ (53)
@@
Under the conditions of Theorem 3.11, any q that the attackers can reproduce is such that $\textbf{q}\notin\mathcal{R}$, i.e.  $\mathcal{R}\cap\mathcal{A}=\emptyset$, and thus
@@
Let $\textbf{q}=(q_{\textsc{c}},q_{\perp},q_{\textsc{i}})$ be the probability distribution corresponding to an attack from Alice and Bob and define $\mathcal{A}$ as the set of all probabilities If they are allowed to answer with probability $\eta\pm\epsilon$, then
$\textbf{q}=(q_{\textsc{c}},1-(\eta+\mu),q_{\textsc{i}}),\mu\in[-\epsilon,\epsilon].$ (55)
@@
$q_{\textsc{i}}\geq\mathfrak{b}_{\eta},\forall\mu\in[-\epsilon,\epsilon].$ (56)
@@
$\mathcal{R}^{0}_{\eta}=\{(p_{\textsc{c}},p_{\perp},p_{\textsc{i}})\in\Delta_{2}\mid p_{\perp}\in[1-(\eta-\epsilon),1-(\eta+\epsilon)]\wedge p_{\textsc{i}}<\mathfrak{b}_{\eta}\},$ (57)
@@
Notice that the strategies $\mathbf{S}_{\textsc{tfkw}}$ and $\mathbf{S}_{guess}$, reproduce $\textbf{q}_{\textsc{tfkw}}=(\cos^{2}(\frac{\pi}{8}),0,1-\cos^{2}(\frac{\pi}{8}))$ and $\textbf{q}_{guess}=(\frac{1}{2},\frac{1}{2},0)$ and, respectively, and therefore $\textbf{q}_{\textsc{tfkw}},\textbf{q}_{guess}\in\mathcal{A}$. The attackers can also do the strategies consisting on always responding $\perp$ or always being incorrect (e.g. answering inconsistent answers $a\neq b$), which reproduce the probabilities $\textbf{q}=(0,1,0)$ and $\textbf{q}=(0,0,1)$, respectively. Therefore, the convex hull of these four points is a convex subset of $\mathcal{A}$. Consider the straight line $s_{1}$ defined by the two points $\textbf{q}_{guess}$ and $(\eta-\epsilon-\mathfrak{b}_{\eta},1-(\eta-\epsilon),\mathfrak{b}_{\eta})\in\partial\mathcal{R}^{0}_{\eta}$, where $\partial$ denotes the boundary, which is given by
$s1:\frac{x-\frac{1}{2}}{\eta-\epsilon-\frac{1}{2}-\mathfrak{b}_{\eta}}=\frac{y-\frac{1}{2}}{\frac{1}{2}-(\eta-\epsilon)}=\frac{z}{\mathfrak{b}_{\eta}}.$ (59)
Consider the set $\mathcal{R}^{1}_{\eta}$, defined as the points in $\Delta_{2}$ that do not belong to $\mathcal{R}^{0}_{\eta}$ that are ’below’ the straight line $s_{1}$ and are such that $p_{\perp}\leq 1-(\eta+\epsilon)$, formally defined as
$\begin{split}\mathcal{R}^{1}_{\eta}=\big{\{}(p_{\textsc{c}},p_{\perp},p_{\textsc{i}})\in\Delta_{2}\mid p_{\textsc{c}}>1-p_{\perp}+\frac{\mathfrak{b}_{\eta}(1-2p_{\perp})}{1-2(\eta-\epsilon)}\wedge p_{\textsc{i}}<\frac{\mathfrak{b}_{\eta}(1-2p_{\perp})}{2(\eta-\epsilon)-1}\big{\}}\setminus\mathcal{R}^{0}_{\eta}.\end{split}$ (60)
@@
Similarly, consider the set $\mathcal{R}^{2}_{\eta}$ consisting of the points in $\Delta_{2}$ that are such that $p_{\perp}\geq 1-(\eta+\epsilon)$ and that are ’below’ the straight line given by the points $\textbf{q}_{\textsc{tfkw}}$ and $(\eta-\epsilon-\mathfrak{b}_{\eta},1-(\eta-\epsilon),\mathfrak{b}_{\eta})$, formally defined as
$\begin{split}\mathcal{R}^{2}_{\eta}=\big{\{}(p_{\textsc{c}},p_{\perp},p_{\textsc{i}})\in\Delta_{2}\mid p_{\textsc{c}}>\frac{\eta-\epsilon-\cos^{2}(\frac{\pi}{8})-\mathfrak{b}_{\eta}}{1-(\eta-\epsilon)}p_{\perp}+\cos^{2}(\frac{\pi}{8})\}\wedge p_{\textsc{i}}<\frac{\mathfrak{b}_{\eta}-\sin^{2}(\frac{\pi}{8})}{1-(\eta-\epsilon)}p_{\perp}+\sin^{2}(\frac{\pi}{8})\big{\}}\end{split}$
@@
Let $\varepsilon_{\textrm{h}}>0$, and $\eta$ and $p_{err}$ be such that $\alpha^{ent}>0$. Then, any sequential strategy that attackers who pre-share $q\leq n/2-5$ qubits to break $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$  is such that $\mathbb{E}[\Gamma_{r}^{ent,att}]\leq 0$ and, moreover, the probability that they are accepted in the $\mathsf{T}^{rf-\mathrm{BB84}}_{\varepsilon_{\textrm{h}}}$ test is exponentially small:
$\mathbb{P}[\Gamma_{r}^{ent,att}\geq r(\alpha^{ent}-\delta)]\leq e^{-r(\alpha^{ent}-\delta)^{2}/2}.$ (61)
Theorem 3.13 shows that there is a way (statistical test) to distinguish an honest prover from attackers with exponentially high probability   i.e.  that after $r$ rounds the attackers will be caught. The values for which $\alpha^{ent}>0$ correspond to the points below the straight line in Fig. 5—the above $\alpha^{ent}-\delta>0$ corresponds to a shift, where $\delta$ can be made small by increasing the number of repetitions.
@@
Let $\textbf{p}_{1}=(959/1000,0,41/1000)$ and $\textbf{p}_{2}=(27/50,23/50,0)$. The line segment given by $\textbf{p}_{1}$ and $\textbf{p}_{2}$ is a subset of $\mathcal{R}$ and it can be described as the intersection of $\Delta_{2}$ and the plane $\gamma_{\textsc{c}}p_{\textsc{c}}-\gamma_{\perp}p_{\perp}-\gamma_{\textsc{i}}p_{\textsc{i}}=0$, with the normal vector $\bm{\gamma}=(\gamma_{\textsc{c}},\gamma_{\perp},\gamma_{\textsc{i}})$, where $\gamma_{\textsc{c}}=\frac{943}{\sqrt{488625947}}\simeq 0.04266$, $\gamma_{\perp}=\frac{1107}{\sqrt{488625947}}\simeq 0.050079$ and $\gamma_{\textsc{c}}=\frac{22057}{\sqrt{488625947}}\simeq 0.99783$. Consider the follow non-empty partition of the simplex $\Delta_{2}$:
@@ 4.1 Discrete uniform choice of basis over the Bloch sphere
Let $\eta$ denote the transmission rate of the qubits sent from $V_{0}$ to P, and let $m_{\theta\varphi}$ be as described above. We define one round of the $\mathrm{QPV}_{m_{\theta\varphi}}^{\eta}$ protocol as follows:
@@
    $V_{0}$ and $V_{1}$ secretly agree on $v\in\{0,1\}$ and $x\in[k]$. Then, $V_{0}$ prepares the qubit state $\left|v_{x}\right\rangle$ as in (69).
@@
    $V_{0}$ sends the qubit $\left|v_{x}\right\rangle$ to P, and $V_{1}$ sends $x$ to P coordinating their times so that they arrive to P at the same time.
@@
    Immediately, P measures the received qubit in the basis $\{\left|0_{x}\right\rangle,\left|1_{x}\right\rangle\}$ and broadcasts her outcome to both verifiers. If, due to experimental losses, the qubit did not arrive, P broadcasts ‘No photon’ with the symbol $\perp$. Denote the response by $v_{\textsf{P}}$.
@@
    If
    1. (a)
      $V_{0}$ and $V_{1}$ receive their respective answers at the time corresponding with the claimed location, and they are equal, i.e.  both receive the same $v_{\textsf{P}}$, then, if
      • $v_{\textsf{P}}=v$, the verifiers record ‘Correct’, denoted by ‘c’,
      • $v_{\textsf{P}}=1-v$, the verifiers record ‘Wrong’, denoted by ‘w’,
      • $v_{\textsf{P}}=\perp$, the verifiers record ‘No photon’, denoted by $\perp$’,
    2. (b)
[…]
At the end of the protocol, when sequentially run $r$ times, the verifiers check that the answers satisfy (1). Notice that $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ is recovered for $m=2$, with the unique choice of $m_{\theta}=2$ and $m_{\varphi}=1$. Analogous to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, for each choice of $m_{\theta\varphi}$, from the analysis of the correlations shown below in Section 4, one can construct a decision test to make the binary decision to either accept or reject, see proof of Theorem 3.13 for how to construct it.
4.2 Security of the $\mathrm{QPV}_{m_{\theta\varphi}}^{\eta}$ in the No-PE model
Analogous to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ in Section 2, for the security analysis, we will consider the purified version of $\mathrm{QPV}_{m_{\theta\varphi}}^{\eta}$, where, in Definition 4.1, in step 1, $V_{0}$ instead of preparing the qubit $\left|0_{x}\right\rangle$, prepares the EPR pair $\left|\Phi^{+}\right\rangle$, sends one qubit register to $P$, and keeps the other register. In a later moment, $V_{0}$ measures his local register in the basis $\{\left|0_{x}\right\rangle,\left|1_{x}\right\rangle\}$. In an analogy to the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol, an attack on the $\mathrm{QPV}_{m_{\theta\varphi}}^{\eta}$ protocol can be associated with a MoE (monogamy-of-entanglement) game in the following way. Let $V$ be the register of the qubit of the verifier, with associated Hilbert space $\mathcal{H}_{V}=\mathbb{C}^{2}$, with $\mathcal{X}=[m]$ and $\mathcal{V}=\{0,1\}$. The verifiers perform the collection of measurements